# uDistil-Whisper: Label-Free Data Filtering for Knowledge Distillation via Large-Scale Pseudo Labelling

### v2

@@ Abstract
Recent work on distilling Whisper’s knowledge into small models using pseudo-labels shows promising performance while reducing the size by up to 50%. This results in small, efficient, and dedicated models. However, a critical step of distillation from pseudo-labels involves filtering high-quality predictions and using only those during training. This step requires ground truth to compare and filter bad examples making the whole process supervised. In addition to that, the distillation process requires a large amount of data thereby limiting the ability to distil models in low-resource settings. To address this challenge, we propose an unsupervised or label-free framework for distillation, thus eliminating the requirement for labeled data altogether. Through experimentation, we show that our best distilled models outperform the teacher model by 5-7 points in terms of WER. Additionally, our models are on par with or better than similar supervised data filtering setup. When we scale the data, our models significantly outperform all zero-shot and supervised models. We demonstrate that it is possible to distill large Whisper models into relatively small models without using any labeled data. Our distilled models are 25-50% more compute and memory efficient while maintaining performance equal to or better than the teacher model.
@@ 1 Introduction
In an effort to democratize Automatic Speech Recognition (ASR), significant attention has been given to multilingual models Pratap et al. (2023b); Radford et al. (2023); Communication et al. (2023a); Zhang et al. (2023). These powerful systems, thanks to their massive number of parameters and training data, can simultaneously transcribe in hundreds of languages. That being said, low-resource languages tend to trail behind in performance compared to counterparts like English. For instance, while OpenAI’s latest ASR model Whisper-large-v3 shows a performance in the single digits on English and Spanish, East Asian and African languages show a much lower performance [^1] https://github.com/openai/whisper/discussions/1762. While the performance reported on these multilingual models is not as low for Arabic, it is important to highlight that it does not fully and accurately represent the diversity of the different varieties that characterize the Arabic language family Abdul-Mageed et al. (2020); Habash (2022). In addition to the disparity in performance, multilingual models are extremely compute-hungry and are thus not equally accessible to everyone. In order to alleviate this issue, knowledge distillation has been the method of choice for multiple works Hinton et al. (2015); Yang et al. (2023); Sanh et al. (2019); Frantar et al. (2022). It has proven to be powerful, effectively reducing model size (and hence compute and memory requirements) while keeping performance on par with the model originally distilled from. Gandhi et al. (2023a) and Waheed et al. (2024) showcase this for English and Arabic, respectively. That being said, one important limitation of these works is their reliance on a large number of high-quality pseudo labels from the teacher model. Ground-truth labels, which are scarce for low-resource languages, are needed to filter out incorrect teacher predictions and ensure such quality. This constraint thus calls for an unsupervised approach to knowledge distillation.
@@
  • We explore various unsupervised methods to filter-out low-quality pseudo-labels, eliminating the necessity of labeled data for distillation.
  • We assess the performance of our distilled models across diverse settings and extend the existing evaluation to include the SADA dataset.
  • We analyze the effectiveness of these metrics in detecting low-quality pseudo-labeled examples.
We organize the paper as follows: we present a review of the existing literature in Section 2 before introducing knowledge distillation along with our methodology in Section 3. Section 4 contains particulars of our experiments and Section 5 delves into the results we obtained. We discuss these results in Section 5, and go over our limitations and conclude our work in Sections 7 and 6, respectively.
@@ 2 Related Work
Knowledge distillation is the process through which knowledge from a trained large model referred to as the teacher is transferred to a smaller model called the student, in order to minimize the requirements at both the compute and memory levels Gandhi et al. (2023a); Gou et al. (2021); Lopes et al. (2017a). Different approaches for knowledge distillation have been explored such as Patient Knowledge Distillation (PKD) Sun et al. (2019), T-S learning Manohar et al. (2018), Cross-Modal Hashing (CMH) Hu et al. (2020) joint Unsupervised Domain Adaptation with KD Zhang et al. (2021) and Data-Free KD Lopes et al. (2017b) among others in the NLP and CV fields alike. Various works have showcased its use specifically for speech recognition, such as Shao et al. (2023) who shrink a Whisper Radford et al. (2023) model tangibly (around 80% of its original size), all while improving its performance, and a Hubert model is reduced by Chang et al. (2021) to 75% of its original size through layer-wise distillation without a significant drop in performance.
Thanks to the rise of multilingual models such as Whisper Radford et al. (2023), Massively Multilingual Speech (MMS) Pratap et al. (2023b), and Seamless-M4T Communication et al. (2023b), low-resource ASR is garnering more attention. Additionally, pseudo-label knowledge distillation is showing promise in the low-resource domain by leveraging these multilingual models to create tailored students that are superior in performance. That being said, the issue of data scarcity becomes more pressing with the reliance of this distillation approach on a large number of ground-truth transcriptions, effectively creating a bottleneck in the pipeline. Doutre et al. (2021) showcase an approach to unsupervised knowledge distillation where they use non-streaming models to distill knowledge into streaming models (which are known to perform worse). Inspired by Liao et al. (2013), they use randomly-segmented audio from YouTube videos to generate pseudo-labels from a non-streaming teacher model. These labels are then used to train the student streaming model. This technique resulted in an improvement in absolute WER between 1.5% and 9.5%. It is important to note, however, that the languages they use (Spanish, French, Italian and Portuguese) are not considered low-resource. Xu et al. (2022) present two novel techniques of self-distillation where the student model is itself used to leverage its own information better. The first technique is dubbed Neighboring Feature Self-Distillation (NSFD) where for every two neighboring layers, the deeper one is distilled into the shallow one, while the second technique, called Attention-based Feature Self-Distillation (AFSD) introduces an attention mechanism that calculates similarity weights to fuse deep features. These methods achieve a reduction of 6.0% and 8.3% in relative CER, respectively. For Arabic speech recognition, Waheed et al. (2024) investigate pseudo-label distillation methods Gandhi et al. (2023a) across various Modern Standard Arabic (MSA) and dialectal datasets. Their exploration demonstrated the efficacy of distillation in enhancing both efficiency and performance, showcasing that smaller dedicated models can outperform larger multilingual ones. While the pseudo-labeling approach has proven effective across diverse languages Gandhi et al. (2023a); Waheed et al. (2024), it traditionally requires labeled data for filtering low-quality pseudo-labeled examples. In our work we address this limitation by introducing unsupervised data filtering methods, thereby eliminating the dependency on labeled data.
@@ 3.1 Knowledge Distillation
Knowledge distillation is the process through which a small student model learns the behavior of a bigger teacher model Hinton et al. (2015); Sanh et al. (2020); Kim and Rush (2016). Gandhi et al. (2023b) introduce pseudo-label distillation, which generates (English) predictions from a teacher Whisper model and filters them through a WER threshold to only keep the most accurate ones. These pseudo-labels are then used to train a smaller student model. Waheed et al. (2024) build on this approach and show that a Whisper student model beats its teacher model in its average performance in an Arabic multi-dialectal setting. We follow this standard student-teacher framework to distill Whisper into small yet powerful models. The objective of the distillation process can be stated as:
@@
  • $\mathcal{L}_{KL}$ represents the Kullback-Leibler (KL) divergence loss, measuring the difference between the distributions of the teacher and student models. It encourages the student model to mimic the probability distribution of the teacher model.
  • $\mathcal{L}_{PL}$ is the objective to train the student model with pseudo labels as ground truth.
The coefficients $\alpha_{KL}$ (0.8) and $\alpha_{PL}$ (1.0) adjust the relative importance of each loss component in the overall knowledge distillation loss $\mathcal{L}_{KD}$.
3.2 Data Filtering
We employ various methods to filter data without relying on ground truth. These methods are based on the teacher model’s logits, synthetic speech for pseudo-labels, proxy references, and multimodal embeddings for speech and pseudo-labels. We describe each method subsequently.
@@ 3.2.1 Proxy Models
This method utilizes a separate, pre-trained ASR model (referred to as proxy model) to generate reference transcripts for the input speech. The quality of the teacher model’s pseudo-label is then assessed based on the Word Error Rate (pWER) between the proxy model’s reference and the teacher’s reference. Lower pWER values indicate better agreement between the teacher and proxy models, suggesting higher confidence and reliability in the pseudo-label. A pre-defined WER threshold can be used to filter out examples exceeding the acceptable error rate. In our work, we use SeamlessM4T-large-v2 as our choice of proxy model.
@@ 3.3 Uncertainty Quantification
We attempt to leverage the uncertainty in the teacher’s output to filter out low-quality examples. Measures derived from the teacher model’s confidence scores are used to identify pseudo-labels with high reliability. Two common uncertainty metrics are employed:
Entropy. Entropy measures the information uncertainty associated with the teacher’s predicted probability distribution over the vocabulary. High entropy indicates the model is unsure about the correct word, suggesting a lower confidence in the pseudo-label. The entropy $H$ for a single utterance with $N$ words is calculated as:
@@
where $p_{i}$ represents the teacher model’s predicted probability for the $i^{\text{th}}$ word in the utterance.
Geometric Mean of Confidence Scores. This metric calculates the geometric mean of individual confidence scores assigned by the teacher model for each word in the pseudo-label. A high geometric mean suggests that the model is confident in all the predicted words, implying a more reliable pseudo-label. The geometric mean $G$ of confidence scores for an $N$-word utterance is:
@@
where $c_{i}$ denotes the confidence score assigned by the teacher model for the $i^{\text{th}}$ word. During filtering, examples with low entropy and a high geometric mean are considered reliable and are retained for training the student model.
3.3.1 Negative Log Likelihood
We make use of an AceGPT-7B Huang et al. (2024) Language Model to compute the log-likelihood of the teacher’s output. A lower negative log-likelihood indicates higher compatibility between the pseudo-label and the language model’s understanding of natural language, suggesting a more accurate pseudo-label. The negative log-likelihood $NLL$ for a teacher-predicted pseudo-label sequence $y$ of length $T$ is:
@@ 3.3.1 Negative Log Likelihood
where $p(y_{t}|y_{1...t-1})$ is the probability assigned by the language model to $y_{t}$, the $t^{\text{th}}$ word in the pseudo-label sequence, given the previous words $(y_{1...t-1})$. Pseudo labels with lower negative log-likelihood values are preferred, as they are more likely to form coherent sequences in the language model, indicating higher quality pseudo labels suitable for training the student models.
@@ 3.3.2 Multimodal Embeddings
In this approach, SONAR Duquenne et al. (2023) is utilized to generate embeddings for both input speech segments and their corresponding pseudo-labels. These embeddings capture the contextual information of the speech signals and the text representations derived from the pseudo-labels. The similarity between these embeddings is computed using the dot product, serving as a measure of how closely aligned the speech and pseudo-label representations are. A high similarity score indicates strong contextual alignment between the speech and text, suggesting that the pseudo-label accurately reflects the content of the speech segment. This helps in identifying and retaining high-quality pseudo-labels while filtering out those that do not adequately represent the speech content. By leveraging the similarity scores derived from SONAR embeddings, we effectively enhance the quality of the pseudo-labels used in training, improving the overall performance as a result. This method enables more accurate model training in scenarios where labeled data is limited or unavailable, optimizing the efficiency and effectiveness of the learning process.
@@ 3.3.3 Similarity with Synthetic Speech
In this approach, we start by generating synthetic speech from the pseudo-labels’ text using XTTS-v2222https://huggingface.co/coqui/XTTS-v2. Next, we compute the similarity between the synthetic speech and the ground truth speech. To assess the quality of the synthetic speech compared to the original speech segment, we utilize metrics such as Perceptual Evaluation of Speech Quality (PESQ) and Mel-Cepstral Distortion (MCD). A high similarity score suggests that the pseudo-labels are likely to have minimal errors. When evaluating these metrics in relation to the Word Error Rate (WER), we observe that PESQ shows a stronger correlation. Therefore, we select PESQ as the primary metric for our experiments. This choice ensures an effective assessment of synthetic speech quality and alignment with the original speech, enhancing the reliability of our pseudo-labeling approach.
@@ 3.5 Teacher and Student Models
In line with previous work Gandhi et al. (2023a); Waheed et al. (2024), we utilize the whisper-large-v2 checkpoint for both pseudo-labeling and as the teacher model during training. We train two variants of students, differing in the number of layers removed from the teacher model. Following (Gandhi et al., 2023b; Waheed et al., 2024), we initialize the student models with maximally spaced layers in the encoder and decoder block of the teacher model. We refer to the models with 16-16 and 32-16 encoder-decoder blocks experimented with 100K segments as UDW-16-16 and UDW-32-16, respectively. The models trained with 500K segments are referred to as UDW-16-16++ and UDW-32-16++.
@@ 4 Experiments
For a fair comparison and analysis, we keep our experimental setup quite identical to Waheed et al. (2024). Apart from evaluating our models on five standard benchmark datasets, and novel dialectal data introduced in Waheed et al. (2024), we extend the evaluation to include the SADA Alharbi et al. (2024) dataset. We describe these corpora below.
4.1 Dataset
FLEURS. The Few-shot Learning Evaluation of Universal Representations of Speech benchmark (FLEURS) Conneau et al. (2022) is an n-way parallel multi-lingual speech dataset and an extension of the FLoRes-101 benchmark Goyal et al. (2022). It comprises 102 languages, each represented by approximately 12 hours of speech. FLEURS is widely utilized for tasks such as ASR, language identification, and speech-to-speech or speech-to-text translation. We use the Arabic split which has MSA speech delivered by speakers with Egyptian accents.
Common Voice. Common Voice (CV) Ardila et al. (2020) is another multilingual dataset powered by volunteer translators and speakers. CV’s latest release to date (CV17.0) counts 124 languages and a total of 31,176 hours. Its Arabic split includes 156 hours of mainly MSA speech.
Multi-Genre Broadcast. Multi-Genre Broadcast (MGB) Ali et al. (2019a, 2017, b) is an Arabic dataset released under three different editions: MGB2 contains 1,200 hours of speech collected from the TV channel Aljazeera Arabic. It is majoritarily MSA speech (> 70%) and is categorized as interviews, reports, and conversations. MGB3, on the other hand, is entirely dialectal with 6 hours of Egyptian speech collected from various YouTube channels. The last edition, MGB5, contains 14 hours of speech extracted from YouTube in the Moroccan dialect.
SADA. The Saudi Audio Dataset for Arabic (SADA) Alharbi et al. (2024) is an Arabic dataset comprised of 668 hours of speech (435 of which are labeled) in both spontaneous and read styles. Most of the speech is in different Saudi dialects (Najdi, Hijazi and Khaleeji among others), but it also includes MSA and dialects from other regions, namely the Levant, Yemen, and Egypt. We report the statistics about the SADA data in Table 1
Dialect Test(S/D) Valid (S/D)
Najdi 1703/2.0709 2249/3.3155
MTOS 1320/4.8044 1048/3.82
Khaliji 1150/1.1308 679/0.6317
Hijazi 809/1.1202 528/0.6423
Unknown 762/0.8325 489/0.4861
NA 167/0.1341 2/0.0004
MSA 157/0.5406 54/0.1682
Egyptian 96/0.0865 45/0.0524
Shamali 18/0.0243 -
Yemeni 7/0.0052 23/0.0349
Levantine - 19/0.0137
Total 6189/10.75 5136/9.17

### v3

# uDistil-Whisper: Label-Free Data Filtering for Knowledge Distillation in Low-Data Regimes

@@ Abstract
Recent work on distilling Whisper’s knowledge into small models using pseudo-labels shows promising performance while reducing the size by up to 50%. This results in small, efficient, and dedicated models. However, a critical step of distillation from pseudo-labels involves filtering high-quality predictions and using only those during training. This step requires ground truth labels to compare and filter low-quality examples making the whole process supervised. In addition to that, the distillation process requires a large amount of data thereby limiting the ability to distill models in low-resource settings. To address this challenge, we propose a distillation framework that does not require any labeled data. Through experimentation, we show that our best distilled models outperform the teacher model by 5-7 points in terms of WER compared to those without filtering and are on par with or perform better than similar supervised data filtering setups. When we scale the data, our models significantly outperform all zero-shot and supervised models. We demonstrate that it is possible to distill large Whisper models into relatively small ones without using any labeled data. Our distilled models are also 25-50% more compute- and memory-efficient while maintaining performance equal to or better than that of the teacher model.
@@ 1 Introduction
In an effort to democratize automatic speech recognition (ASR), significant attention has been given to multilingual models Pratap et al. (2023); Radford et al. (2023); Communication et al. (2023); Zhang et al. (2023). These powerful systems, thanks to their massive number of parameters and training data, can simultaneously transcribe in hundreds of languages. That being said, low-resource languages tend to trail behind in performance compared to high-resource counterparts such as English Radford et al. (2023). For instance, while OpenAI’s latest ASR model Whisper-large-v3 shows the word error rate (WER) in the single digits on English and Spanish, a much lower performance is shown for East Asian and African languages [^1] https://github.com/openai/whisper/discussions/1762 Talafha et al. (2023). While the performance reported on these multilingual models is not as low for Arabic, it is important to highlight that it does not fully and accurately represent the diversity of the different varieties that characterize the Arabic language family Abdul-Mageed et al. (2020); Habash (2022). In addition to the disparity in performance, multilingual models are extremely compute-intensive and are thus not equally accessible to everyone. In order to alleviate this issue, knowledge distillation has been the method of choice for multiple works Hinton et al. (2015); Yang et al. (2023); Sanh et al. (2020); Frantar et al. (2022).
Knowledge distillation has proven to be highly effective in reducing model size, thereby lowering compute and memory requirements while maintaining performance comparable to larger teacher models. For example, Gandhi et al. (2023) and Waheed et al. (2024) demonstrate its effectiveness for English and Arabic, respectively. However, a significant limitation in these works is their reliance on ground-truth labels to filter out low-quality pseudo-labels generated by the teacher model—a resource that is often scarce, particularly for low-resource languages. Additionally, Waheed et al. (2024) shows that training on unfiltered data leads to suboptimal performance. This reliance on labeled data underscores the need for an unsupervised approach to data filtering in knowledge distillation, enabling distilled models to perform better without depending on any ground-truth labels.
@@
  • We explore various unsupervised methods to filter out low-quality pseudo-labels, eliminating the need of labeled data for distillation.
  • We assess the performance of our distilled models across diverse settings and extend the existing evaluation to include new datasets.
  • We analyze the effectiveness of our best metrics in detecting low-quality pseudo-labeled examples.
We organize the paper as follows: we present a review of the existing literature in Section 2 before introducing knowledge distillation along with our methodology in Section 3. Section 4 contains particulars of our experiments and Section 5 delves into the results we obtain. We discuss these results in Section 5, conclude in Section 6, and outline our work’s limitations in Section 7.
@@ 2 Related Work
Scaling speech foundation models in both data and model size has led to systems capable of handling a wide range of speech tasks Pratap et al. (2023); Zhang et al. (2023); Radford et al. (2023); Communication et al. (2023). These models effectively transcend language barriers, demonstrating robust performance across numerous languages, including many low-resource ones Hsu et al. (2024); Yeo et al. (2024). However, these models face two key limitations: (1) their large size Radford et al. (2023); Communication et al. (2023), which results in high computational costs, and (2) the under-representation of low-resource languages in training, leading to suboptimal performance on these languages Radford et al. (2023); Talafha et al. (2023). To address the first issue various efficient decoding methods SYSTRAN ; Segal-Feldman et al. (2024); Malard et al. (2023); Leviathan et al. (2023) were proposed. However, resulting models do not necessarily enjoy a reduced memory footprint.
@@
To address the efficiency issue, knowledge distillation Hinton et al. (2015); Gou et al. (2021) is proven to be very effective. Different approaches for knowledge distillation have been explored such as Patient Knowledge Distillation (PKD) Sun et al. (2019), T-S learning Manohar et al. (2018), Cross-Modal Hashing (CMH) Hu et al. (2020), Joint Unsupervised Domain Adaptation with KD Zhang et al. (2021) and Data-Free KD Lopes et al. (2017) among others in the NLP and CV fields alike. Knowledge distillation has also been applied to speech recognition:  Shao et al. (2023) shrink a Whisper Radford et al. (2023) model tangibly ($\sim 80\%$ of its original size), all while improving its performance;  Chang et al. (2021) reduce a Hubert model by to 75% of its initial size through layer-wise distillation without a significant drop in performance.
Knowledge distillation has also been applied to speech-to-text tasks Nayem et al. (2023); Hentschel et al. (2024); Tian et al. (2022) to reduce the memory and compute requirements. More recently, it has been used to distill multilingual models from Whisper Ferraz et al. (2024) and strong monolingual models Gandhi et al. (2023), as well as in low-resource settings. An example of the latter is Waheed et al. (2024) who investigate pseudo-label distillation methods Gandhi et al. (2023) across various Modern Standard Arabic (MSA) and dialectal Arabic datasets. Their exploration demonstrate the efficacy of distillation in enhancing both efficiency and performance, showcasing that smaller dedicated models can outperform larger multilingual ones. While the pseudo-labeling approach has proven effective across diverse languages Gandhi et al. (2023); Waheed et al. (2024), it traditionally requires labeled data for filtering low-quality pseudo-labeled examples. In this work we address this limitation by introducing an unsupervised framework based on data filtering methods, thereby eliminating the dependency on labeled data altogether.
@@ 3.1 Knowledge Distillation
Knowledge distillation is a framework through which a small student model learns the behavior of a bigger teacher model Hinton et al. (2015); Sanh et al. (2020); Kim and Rush (2016). Gandhi et al. (2023) introduce knowledge distillation via pseudo-labeling, which generates (English) predictions from a teacher Whisper model and filters them through a WER threshold to only keep the most accurate ones. These pseudo-labels are then used to train a smaller student model. Waheed et al. (2024) build on this approach and show that a Whisper student model beats its teacher model in its average performance in an Arabic multi-dialectal setting. We follow this standard student-teacher framework to distill Whisper into small yet powerful models. The objective of the distillation process can be stated as:
@@
  • $\mathcal{L}_{KL}$ is the Kullback-Leibler (KL) divergence loss, encouraging the student model to match the teacher’s probability distribution.
  • $\mathcal{L}_{PL}$ trains the student using pseudo labels as ground truth.
The coefficients $\alpha_{KL}$ (0.8) and $\alpha_{PL}$ (1.0) balance the contributions of each loss in the overall distillation loss $\mathcal{L}_{KD}$.
3.2 Label-Free Data Filtering
Previous data filtering methods, such as those used by Gandhi et al. (2023) and Waheed et al. (2024), rely on ground-truth labels by calculating WER between teacher-generated pseudo-labels and reference transcripts, discarding examples with high WER. However, this dependence on labeled data limits their use in low-resource languages where such labels are scarce.
Our approach introduces label-free filtering methods that assess pseudo-label quality using the teacher model’s logits, synthetic speech, proxy models, and multimodal embeddings. Each method is detailed below.
@@ 3.2.1 Proxy Models
We use a trained ASR model, SeamlessM4T-large-v2 Communication et al. (2023), as a proxy to generate reference transcripts for the input speech. Then, the quality of pseudo-labels generated by the teacher model is evaluated by calculating the WER between the proxy’s and the teacher’s outputs, referred to as pWER. Lower pWER values indicate higher agreement and examples exceeding a defined pWER threshold are discarded.
@@ 3.3 Uncertainty Quantification
We leverage uncertainty in the teacher’s output to filter low-quality examples, using two common metrics:
Entropy. Entropy measures uncertainty in the teacher’s predicted probability distribution. High entropy suggests low confidence. It is calculated as:
@@
where $p_{i}$ is the predicted probability for the $i^{\text{th}}$ word.
Geometric Mean of Confidence Scores. The geometric mean of the confidence scores, representing the probability of each decoded token in the teacher-generated pseudo-label, is used to assess the overall confidence in the pseudo-label. It is calculated as:
@@
where $c_{i}$ is the confidence score for each word. Low entropy and high geometric mean indicate better pseudo-labels.
3.3.1 Negative Log-Likelihood
We compute the negative log-likelihood (NLL) using AceGPT-7B Huang et al. (2024) to assess the quality of pseudo-labels. Lower NLL suggests higher pseudo-label accuracy. It is calculated as:
@@ 3.3.1 Negative Log-Likelihood
where $p(y_{t}|y_{1...t-1})$ is the probability assigned to the $t^{\text{th}}$ word.
@@ 3.3.2 Multimodal Embeddings
We use SONAR Duquenne et al. (2023) to generate embeddings for input speech and pseudo-labels. The dot product is used to compute similarity, with higher scores indicating better alignment and pseudo-label quality.
@@ 3.3.3 Similarity with Synthetic Speech
Synthetic speech is generated from pseudo-labels using XTTS-v2222https://huggingface.co/coqui/XTTS-v2 and compared with input speech. Similarity is assessed via PESQ and MCD, where high PESQ and low MCD scores indicate higher label accuracy.
@@ 3.5 Teacher and Student Models
In line with previous work Gandhi et al. (2023); Waheed et al. (2024), we utilize the whisper-large-v2 checkpoint [^3] https://huggingface.co/openai/whisper-large-v2 for both pseudo-labeling and as the teacher model during training. We train two variants of students, differing in the number of layers removed from the teacher model. Following Gandhi et al. (2023) and Waheed et al. (2024), we initialize the student models with maximally spaced layers in the encoder and decoder block of the teacher model. We refer to the models with 16-16 and 32-16 encoder-decoder blocks as UDW-16-16 and UDW-32-16, respectively, when we train with 100K segments and as UDW-16-16++ and UDW-32-16++ when we train with 500K segments.
@@ 4 Experiments
For a fair comparison and analysis, we keep our experimental setup quite identical to Waheed et al. (2024). Apart from evaluating our models on five standard benchmark datasets, and novel dialectal data introduced in Waheed et al. (2024), we extend the evaluation to include the SADA Alharbi et al. (2024) dataset. We describe our evaluation datasets below.
4.1 Evaluation Dataset
FLEURS. The Few-shot Learning Evaluation of Universal Representations of Speech (FLEURS) Conneau et al. (2022) is a multilingual speech dataset with 102 languages, each with  12 hours of speech. It supports tasks like ASR, language identification, and translation. We use the Arabic subset, which features MSA speech with Egyptian accents.
Common Voice. Common Voice (CV) Ardila et al. (2020) is a volunteer-driven multilingual dataset with 124 languages and 31,176 hours of speech. The Arabic subset includes 156 hours of primarily MSA speech.
Multi-Genre Broadcast. The MGB dataset Ali et al. (2019a, 2017, b) consists of MGB2 (1,200 hours of MSA-dominant Aljazeera Arabic speech), MGB3 (6 hours of Egyptian dialectal speech), and MGB5 (14 hours of Moroccan dialectal speech).
SADA. The Saudi Audio Dataset for Arabic (SADA) Alharbi et al. (2024) includes 668 hours of Arabic speech (435 labeled), featuring Saudi dialects, MSA, and speech from the Levant, Yemen, and Egypt. Statistics are provided in Appendix B 2.1.